The function f(x,y,z)=max{y/x} is a homogeneous function on R².is it true or false.give reasons for your answer


1
Expert's answer
2021-02-03T13:41:02-0500

"Solution:~ False.
\\~~~~~ A ~homogeneous ~ real ~ valued~ function~ of ~two ~variable ~ x ~ and ~y ~ 
\\is ~a ~real ~ valued~ function~ that ~satisfies ~the ~condition~ f(rx,ry)=r^k f(x,y),
\\ for ~some ~constant ~ k~ and ~all ~real ~ number ~ r.
\\~~~Now,~ the ~given ~ real ~valued ~ function ~ of ~ two ~variable~ is ~ f(x,y)=max\{\frac{y}{x},x\}
\\Putting ~ x=rx, ~ and ~ y=ry~ , we ~get,
\\f(rx,ry)=max\{\frac{ry}{rx},rx\}
\\~~~~~~~~~~~~~~~~~=max\{\frac{y}{x},rx\} 
\\~~~~~~~~~~~~~~~~~\neq r^k f(x,y)~ ,for ~some ~constant ~ k~ and ~all ~ real ~ number ~ r.
\\Hence ,f~ is ~not ~ a ~ homogeneous ~function."